Subject: Minutes of APO 3.5m user's committee meeting, 2/8/99

From: strauss@astro.Princeton.EDU

Submitted: Fri, 12 Feb 1999 11:40:18 -0500 (EST)

Message number: 337 (previous: 336, next: 338 up: Index)

	Apache Point Observatory 3.5m User's Committee Meeting
****************February 9, 1998**********************

Attending: Rene Walterbos, Michael Strauss, Bruce Gillespie, Alan
Uomoto, Ed Turner, Jeff Secker, Ed Kibblewhite, Chris Stubbs

3.5m special BoG meeting
Site report
	Recent net performance
	Image quality tests
Secondary status
DDT status
Instrument reports
	DIS, SPIcam, Guider, Echelle, GRIM2

3.5m special BoG meeting
  Last BoG meeting had a discussion of policy and strategic issues
(such as schemes for funding new instruments).  They are interested in
a follow-up meeting or workshop with some subset of the board to
discuss this in more detail, and are interested in having the User's
Committee involved.  The idea is to have a series of recommendations
or policy decisions to be made on the spot, or to be approved in a
future BoG meeting.  One possibility is to have this meeting in June
at APO; alternatively, to have it in August together with a community

Site report
  Gillespie: Network performance has been a big issue lately.  Users
have been complaining a lot about latency, and lots of down-time.  The
problem is that the network services has been continuing to degrade.
No one single bottleneck is apparent.  There are solutions, but they
are all quite expensive.  Craig Loomis will write up a report of the
APO side of things and distribute it.  Uomoto: We need to get a Las
Cruces expert who has experience working with networks out of New
Mexico.  We'll discuss this in detail either by sub-committee, or at
the next meeting, after we've had a chance to read Loomis' paper.  A
related issue is future upgrades of the remote observing software, in
the context of the 3-yr plan.
  One immediate solution:  Connection via modem to the site.  An
action item for Gillespie: to recheck what the modem links are at the
  Stubbs: we should run Loomis' report by some network-savvy people. 

  Stubbs: Jeff Morgan et al have been carrying out tests of a
Shack-Hartmann image-quality device; it takes the measurement of the
wavefront and gives you instructions on how to tune things up (e.g.,
"move the secondary in such a such a way").  It allows you to do
collimation and alignment in real time.  Alignment problems remain; in
addition, high-frequency hash in the PSF is probably due to quilting
errors in the secondary (although this is not for sure).  Jeff Morgan
is recommending that we buy this device (of order $30 K).  There is
only moderate interest among the SDSS people in using such a device;
unless we spend much more money, it would probably not be
interchangeable between the two telescope.

  Gillespie: KPNO has a thermal camera (10-12 micron) to do a thermal
audit of several telescopes.  We've borrowed it, and we've made maps
of the thermal environment around the 3.5m.  Several thermal leaks
were found (some of them were surprises), some of which are easy to
fix (and some are more difficult).  The instruments and the guider
themselves are heat sources; starting to think about how to address

  Gillespie: Regular throughput measurement program was started last
fall.  Charles Corson says the method we've been using is not as
accurate as we would like, due to some issues of extinction, etc.
Gillespie will send around Corson's e-mail with the details.  We
should be able to use the SDSS Monitor Telescope data for measuring
extinction.  In any case, we should rethink the use of DIS in
spectroscopic mode for these measurements; we might simply work off
direct imaging data (from SPICAM?  from the guide camera?).  Do we
want a rough sanity check on the throughput, or do we want a
throughput measurement to 1% to look for subtle long-term changes?  Of
course, we want both, but all this needs to be thought through.
Gillespie will distribute a report from Corson.

  Recruiting another 3.5m observing specialist.  Charles Corson is
leaving us, to take a job with WIYN.

  Secondary: Secondary mirror, polishing cell are at Swales right now.
Things seem to be on schedule.  There was some discussion of looking
for people to make a few outstanding parts; possibilities of doing it
at UW or at Sunspot.  Uomoto: A silver coating, rather than aluminum?
Kibblewhite: If you're not careful, silver coatings can be quite
problematic; bad interaction with dust in the air. 
  Remember, the secondary and tertiary can be coated at Sunspot. 
The question: *if* we can protect these coatings, and *if* there are
not science drivers to keep the UV throughput of the telescope up, it
is a reasonable thing to do. 

  Only one DDT proposal has been received, for gamma-ray burst

  Instrument reports: 
DIS: the hose that feeds the dewars has developed a leak.  Looking
into getting a larger feed dewar, and repairs on the hose.  Gunn and
Lucinio will look at the read noise problem on DIS.  The new
medium-resolution gratings seem to be very good; good throughput; good
resolution.  On red: 70% throughput of low-res grating.  Blue: 90%
throughput of low-res grating.  It is easy to switch these in and out
(although you don't want to do it in excess, and so it requires some
thought on scheduling).  Do we allow grating changes in the middle of
the night?  Only an issue if you need all three sets of gratings in
one night.

SPICAM: Some problems with the shutter sticking.  Also some problems
with the cryogenic system.

Guider: Some problems with the guider Mcintosh, and this computer has
been going down fairly often lately.  But this is the last cleanup of
this system; it has been working very well.  There have been 10-15
minute exposures with 0.8" stars; wow!

Echelle: Still in Chicago.  Things are progressing well, but very
slowly.  Ed Kibblewhite says it will likely go to the mountain in May.
The hope is that commissioning will take of order 3 months.

GRIM: Reports of defective data frames recently.  These reports have
been forwarded on to Mark Hereld; we have not heard back. 

  Previous month's minutes are approved. 

  Next meeting, March 9, 1998, 12:30 PM

APO APO APO APO APO  Apache Point Observatory 3.5m  APO APO APO
APO  This is message 337 in the apo35-general archive. You can find
APO  the archive on
APO  To join/leave the list, send mail to
APO  To post a message, mail it to